OpenStudy (anonymous):

The square of x is equal to 4 time the square of y. If x in one more than twice y, what is the value of x. a(-4) b(-1/2) c(-1/4) d(1/4) e(1/2

OpenStudy (anonymous):

x^2 = 4y^2 x = 2y + 1 Substitute 2y + 1 for x in the first equation so that you only have one variable. (2y+1)^2 = 4y^2 (2y+1)(2y+1) = 4y^2 4y^2 + 4y + 1 = 4y^2 4y + 1 = 0 4y = -1 y = -1/4 x = 2y + 1 x = 2(-1/4) + 1 x = -1/2 + 1 x = -1/2

OpenStudy (anonymous):

so you sub the x into the y. Thanks!